OpenStudy (anonymous):

A flat circular plate of copper has a radius of 0.169 m and a mass of 35.8 kg. What is the thickness of the plate? Answer in units of m.

OpenStudy (anonymous):

The volume of a cylinder: \[(1): v=\pi r ^{2}*h\] where r is the radius and h the height. The density of copper is \[(2): d = 8.94g / cm^{3}\] where density is \[(3): d=m/v\] So now we know that \[(4): 8.94 = 35800/v\] Solve eq. (4) for v and substitute in eq. (1). Then substitute r^2 in eq. 1 for the known radius (in cm) and solve for h. :)

OpenStudy (anonymous):

Of course, that result will be in cm - since we used density in g/cm^3, we need to use r in cm in eq. 1, so after solving for h it'll be cm as well.